# Flexible and efficient spatial extremes emulation via variational autoencoders

### v3

@@ Appendix D XVAE details
In this section, we will illustrate the details of Eqs. (5) and (8). Recall the encoder in the XVAE encodes the information in $\boldsymbol{x}_{t}$, $t=1,\ldots,n_{t}$, using a three-layer perceptron neural network. The three-layer perceptron neural network has the form of:
@@
$\begin{split}\boldsymbol{l}_{1,t}&=\mathrm{relu}(\boldsymbol{W}_{5}\boldsymbol{z}_{t}+\boldsymbol{b}_{5}),\\ \boldsymbol{l}_{2,t}&=\mathrm{relu}(\boldsymbol{W}_{6}\boldsymbol{l}_{1,t}+\boldsymbol{b}_{6}),\\ (\alpha_{t},\boldsymbol{\gamma}_{t}^{T})^{T}&=\mathrm{relu}(\boldsymbol{W}_{7}\boldsymbol{l}_{2,t}+\boldsymbol{b}_{7}),\\ \boldsymbol{y}_{t}&=(\boldsymbol{\Omega}^{1/\alpha_{t}}\boldsymbol{z}_{t})^{\alpha_{0}},\end{split}$ (D.3)
in which $\boldsymbol{\Omega}=(\boldsymbol{w}_{1},\cdots,\boldsymbol{w}_{n_{s}})^{T}$ is a $n_{s}\times K$ matrix with its $j$th row being $\boldsymbol{w}_{j}^{T}=(\omega_{1j},\ldots,\omega_{Kj})$. The weights $\{\boldsymbol{W}_{5},\ldots,\boldsymbol{W}_{7}\}$ and biases $\{\boldsymbol{b}_{5},\ldots,\boldsymbol{b}_{7}\}$ combined are denoted by $\boldsymbol{\phi}_{d}$, in which $\boldsymbol{W}_{5}$ and $\boldsymbol{W}_{6}$ are both $K\times K$ matrices while $\boldsymbol{W}_{7}$ is a $(K+1)\times K$ matrix, and $\boldsymbol{b}_{5}$ and $\boldsymbol{b}_{6}$ are $K\times 1$ vectors while $\boldsymbol{b}_{7}$ is a $(K+1)\times 1$ vector.
D.1 Reparameterization trick
@@
which can be approximated using Monte Carlo as shown in Eq. (2). However, it is not straightforward to approximate the partial derivative of the ELBO with respect to $\boldsymbol{\phi}_{e}$ (denoted by $\nabla_{\boldsymbol{\phi}_{e}}\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}$), which is needed in the stochastic gradient descent algorithm. Since the expectation in ELBO is taken under the distribution $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$.
@@
in which $\boldsymbol{\eta}_{t}=(\eta_{1t},\eta_{2t},\cdots,\eta_{Kt})^{T}$ and $\eta_{kt}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,1)$. As a consequence, the Jacobian of the transformation from $\boldsymbol{Z}_{t}$ to $\boldsymbol{\eta}_{t}$ is
@@
where $\boldsymbol{\eta}^{l}$, $l=1,\ldots,L$, are independent draws from $N(\boldsymbol{0}_{K},\boldsymbol{I}_{K\times K})$ and $\boldsymbol{Z}^{l}=\boldsymbol{\mu}_{t}+\boldsymbol{\zeta}_{t}\odot\boldsymbol{\eta}^{l}$. Also, $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t}\mid\boldsymbol{z}^{l})$ and $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}^{l})$ are defined in Eqs. (7) and (6). Furthermore,
@@
D.2 Effect of knot locations
@@
$x^{*}\leftarrow$ $u$th quantile of the concatenated vector $(\boldsymbol{x}_{1}^{T},\cdots,\boldsymbol{x}_{n_{t}}^{T})^{T}$;
@@
Figures 1(e) and 1(f) compare the simulated and emulated spatial fields of the 50th replicate by plotting their quantiles against each other (when pooling the spatial data into the same plot). We see that both emulations align very well with the simulated data set. Although this might not be the most appropriate way of evaluating the quality of the emulations because there is spatial dependence and non-stationarity within each spatial replicate, QQ-plots still provide value in determining whether the spatial distribution is similar at all quantile levels, which is complementary to the empirical $\chi_{ij}(u)$ described in Section 2.2.
@@
D.3 Stochastic gradient descent optimization
A major advantage of approximating the ELBO as presented in Eq. (2) lies in the ability to perform joint optimization over all parameters ($\boldsymbol{\phi}_{e}$ and $\boldsymbol{\phi}_{d}$) using stochastic gradient descent (SGD). This optimization is efficiently implemented using a tape-based automatic differentiation module called autograd within the R package torch (Falbel and Luraschi, 2023). Built on PyTorch, this package offers rapid array computation, leveraging robust GPU acceleration for enhanced computational efficiency. It stores all the data inputs and VAE parameters in the form of torch tensors, which are similar to R multi-dimensional arrays but are designated for fast and scalable matrix calculations and differentiation.
@@ D.3 Stochastic gradient descent optimization
Importantly, our XVAE algorithm can scale efficiently to massive spatial data sets. The existing max-stable, inverted-max-stable, and other spatial extremes models are limited to applications with less than approximately $1,000$ locations using a full likelihood or Bayesian approach; see Section A for more details on these alternative approaches. By contrast, our approach can fit a globally non-stationary spatial extremes process, with parameters evolving over time, to a data set of unprecedented spatial dimension of more than $16,000$ locations, and also facilitates data emulation in such dimensions. See Section 4 for details.
@@
$(\boldsymbol{\phi}_{e}^{(j)},\boldsymbol{\phi}_{d}^{(j)})^{T}\leftarrow$ Initialized parameters;
@@
             $(\boldsymbol{\mu}_{t}^{T},\log\boldsymbol{\zeta}_{t}^{T})^{T}\leftarrow\mathrm{EncoderNeuralNet}_{\boldsymbol{\phi}_{e}^{(j)}}(\boldsymbol{x}_{t})$;
@@
             $(\alpha_{t},\boldsymbol{\gamma}_{t}^{T})^{T}\leftarrow\mathrm{DecoderNeuralNet}_{\boldsymbol{\phi}_{d}^{(j)}}(\boldsymbol{z}_{t})$;
@@
       Apply update: $(\boldsymbol{\phi}_{e}^{(j+1)},\boldsymbol{\phi}_{d}^{(j+1)})^{T}\leftarrow(\boldsymbol{\phi}_{e}^{(j)},\boldsymbol{\phi}_{d}^{(j)})^{T}+\boldsymbol{v}$;
       $\boldsymbol{L}\leftarrow(\boldsymbol{L}^{T},\mathcal{L}_{\boldsymbol{\phi}_{e}^{(j)},\boldsymbol{\phi}_{d}^{(j)}}(\mathcal{M}))^{T}$ ;
@@
Algorithm 2 Stochastic Gradient Descent with momentum to maximize the ELBO defined in Eq. (2). We set $|\mathcal{M}|=n_{t}$ and $L=1$ in our experiments.
D.4 Finding starting values
@@
$\hat{\boldsymbol{z}}_{t}\approx\{(\boldsymbol{\Omega}^{\frac{1}{\alpha}})^{T}\boldsymbol{\Omega}^{\frac{1}{\alpha}}\}^{-1}(\boldsymbol{\Omega}^{\frac{1}{\alpha}})^{T}\boldsymbol{x}_{t}^{\frac{1}{\alpha_{0}}},\;t=1,\ldots,n_{t}.$
We use QR decomposition to solve the following linear system to get the initial value $\boldsymbol{W}_{1}^{(0)}$: $(\hat{\boldsymbol{z}}_{1},\cdots,\hat{\boldsymbol{z}}_{n_{t}})^{T}=(\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n_{t}})^{T}\boldsymbol{W}_{1}^{T}$. Also, set $\boldsymbol{b}_{1}^{(0)}=(0,\ldots,0)^{T}$. The initial values of $\boldsymbol{h}_{1,t}$ in Eq. (D.1) satisfy $\boldsymbol{h}_{1,t}\approx\hat{\boldsymbol{z}}_{t}$, $t=1,\ldots,n_{t}$.
@@
Appendix E Additional figures from the simulation study
We show additional figures that are complementary to those included in Section 3. Figure E.1 displays the simulated data sets from Models I, II, IV and V and their emulated fields using both XVAE and hetGP. See Figure 2 for comparison for Model III. Figure E.2 displays QQ-plots from the spatial data to compare the overall distributions of the simulated and emulated data sets. Figure E.3 compares the empirically estimated $\chi_{h}(u)$ as described in Section 2.2 from the data replicates simulated from Models I, II, IV and V and their emulations at three different distances $h=0.5,2,5$ under the working assumption of stationarity. Figure E.4 shows the estimates of $\mathrm{ARE}_{\psi}(u)$ defined in Eq. (13), $\psi=0.05$, for both simulations and XVAE emulations under Models I, II, IV and V. See Figure 4 for $\chi_{h}(u)$ and $\mathrm{ARE}_{\psi}(u)$ estimates for Model III. Lastly, Figure E.5 shows coverage probabilities of $\{\gamma_{kt}:k=1,\ldots,K\}$ for $t=1$ from fitting Model III. Coverage probabilities when $\gamma_{k}=0$ are poor, though upper bounds of credible intervals are consistently less than $10^{-6}$.
Figure E.1: Simulated data sets (left column) and emulated fields (XVAE, middle column; hetGP, right column) from Models I, II, IV and V (top to bottom). In all cases, we use data-driven knots for emulation using XVAE. See Figure 2 for comparison for Model III.
@@ Appendix E Additional figures from the simulation study
Figure E.3: The empirically-estimated tail dependence measure $\chi_{h}(u)$ at $h=0.5$ (left), 2 (middle), 5 (right) for Models I, II, IV and V (top to bottom), based on simulated (black) and XVAE emulated (red) data. See Figure 4 for $\chi_{h}(u)$ estimates for Model III.
Figure E.4: Estimates of $\mathrm{ARE}_{\psi}(u)$, $\psi=0.05$, for both simulations (black) and XVAE emulations (red) under Models I, II, IV and V (left to right). See the right panel of Figure 4 for $\mathrm{ARE}_{\psi}(u)$ estimates for Model III.

### v4

@@ Appendix D XVAE details
D.1 General framework
In this section, we will illustrate the details of Eqs. (8) and (11) in the main paper. Recall the encoder in the XVAE encodes the information in $\boldsymbol{x}_{t}$, $t=1,\ldots,n_{t}$, using a three-layer perceptron neural network. The three-layer perceptron neural network has the form of:
@@ D.1 General framework
Here, we opted to not use a heavy-tailed distribution for the variational distribution $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ for a few reasons: (1) the variational distribution serves to “approximate” the true posterior distribution $p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$. Many variational Bayesian methods (see Eq. (4) of Maceda et al., 2024, for example) use a heteroskedastic Gaussian model $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ to approximate $p_{\boldsymbol{\theta}}(\boldsymbol{z}\mid\boldsymbol{x})$. This choice is theoretically supported by an adaptation of Bernstein–von Mises Theorem to the spatial context under a form of spatial mixing condition, which basically requires that observations become “effectively independent” as the distance between them grows—a condition met by our compactly supported Wendland basis functions. Using the terminology outlined in Bradley (2005), we can verify that our max-id model satisfies the more stringent $\phi$-mixing conditions. (2) We experimented with Pareto-tailed variational distributions and they underperformed compared to Gaussian distributions. Intuitively, the mean vector $\boldsymbol{\mu}_{t}$ in $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ anchors the encoding’s center, while the standard deviation $\boldsymbol{\zeta}_{t}$ determines the range of variation around this center. Since the prior distribution $p_{\boldsymbol{\theta}}(\boldsymbol{z})$ is already heavy-tailed, allowing the variational distribution to diverge too widely from the mean proved counterproductive. (3) The variational distribution is regularized by the evidence lower bound (ELBO), in which we try to minimize the KL distance between $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$ and $p_{\boldsymbol{\theta}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$. As long as the ELBO converges, we believe using the heteroskedastic Gaussian model as the variational distribution is sufficient, as evidenced by the extensive simulation results presented in the paper.
@@
$\begin{split}\boldsymbol{l}_{1,t}&=\mathrm{relu}(\boldsymbol{W}_{5}\boldsymbol{z}_{t}+\boldsymbol{b}_{5}),\\ \boldsymbol{l}_{2,t}&=\mathrm{relu}(\boldsymbol{W}_{6}\boldsymbol{l}_{1,t}+\boldsymbol{b}_{6}),\\ (\alpha_{t},\boldsymbol{\gamma}_{t}^{\top})^{\top}&=\mathrm{relu}(\boldsymbol{W}_{7}\boldsymbol{l}_{2,t}+\boldsymbol{b}_{7}),\\ \boldsymbol{y}_{t}&=(\boldsymbol{\Omega}^{1/\alpha_{t}}\boldsymbol{z}_{t})^{\alpha_{0}},\end{split}$ (D.3)
in which $\boldsymbol{\Omega}=(\boldsymbol{w}_{1},\cdots,\boldsymbol{w}_{n_{s}})^{\top}$ is a $n_{s}\times K$ matrix with its $j$th row being $\boldsymbol{w}_{j}^{\top}=(\omega_{1j},\ldots,\omega_{Kj})$. The weights $\{\boldsymbol{W}_{5},\ldots,\boldsymbol{W}_{7}\}$ and biases $\{\boldsymbol{b}_{5},\ldots,\boldsymbol{b}_{7}\}$ combined are denoted by $\boldsymbol{\phi}_{d}$, in which $\boldsymbol{W}_{5}$ and $\boldsymbol{W}_{6}$ are both $K\times K$ matrices while $\boldsymbol{W}_{7}$ is a $(K+1)\times K$ matrix, and $\boldsymbol{b}_{5}$ and $\boldsymbol{b}_{6}$ are $K\times 1$ vectors while $\boldsymbol{b}_{7}$ is a $(K+1)\times 1$ vector.
D.2 ELBO empirical estimates
Therefore, the ELBO can be approximated by Monte Carlo as
$\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x})\approx\frac{1}{L}\sum_{l=1}^{L}\log\frac{p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x},\boldsymbol{Z}^{l})}{q_{\boldsymbol{\phi}_{e}}(\boldsymbol{Z}^{l}\mid\boldsymbol{x})},$ (D.4)
where $\boldsymbol{Z}^{1},\ldots,\boldsymbol{Z}^{L}$ are independent draws from $q_{\boldsymbol{\phi}_{e}}(\cdot\mid\boldsymbol{x})$. If there are replicates of the process, $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n_{t}}$, then $\sum_{t=1}^{n_{t}}\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t})$ is considered.
D.3 Reparameterization trick
@@
which can be approximated using Monte Carlo as shown in Eq. (D.4). However, it is not straightforward to approximate the partial derivative of the ELBO with respect to $\boldsymbol{\phi}_{e}$ (denoted by $\nabla_{\boldsymbol{\phi}_{e}}\mathcal{L}_{\boldsymbol{\phi}_{e},\boldsymbol{\phi}_{d}}$), which is needed in the stochastic gradient descent algorithm. Since the expectation in ELBO is taken under the distribution $q_{\boldsymbol{\phi}_{e}}(\boldsymbol{z}_{t}\mid\boldsymbol{x}_{t})$.
@@
in which $\boldsymbol{\eta}_{t}=(\eta_{1t},\eta_{2t},\cdots,\eta_{Kt})^{\top}$ and $\eta_{kt}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,1)$. As a consequence, the Jacobian of the transformation from $\boldsymbol{Z}_{t}$ to $\boldsymbol{\eta}_{t}$ is
@@
where $\boldsymbol{\eta}^{l}$, $l=1,\ldots,L$, are independent draws from $N(\boldsymbol{0}_{K},\boldsymbol{I}_{K\times K})$ and $\boldsymbol{Z}^{l}=\boldsymbol{\mu}_{t}+\boldsymbol{\zeta}_{t}\odot\boldsymbol{\eta}^{l}$. Also, $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{x}_{t}\mid\boldsymbol{z}^{l})$ and $p_{\boldsymbol{\phi}_{d}}(\boldsymbol{z}^{l})$ are defined in Eqs. (10) and (9) of the main paper. Furthermore,
@@
D.4 Effect of knot locations
@@
$x^{*}\leftarrow$ $u$th quantile of the concatenated vector $(\boldsymbol{x}_{1}^{\top},\cdots,\boldsymbol{x}_{n_{t}}^{\top})^{\top}$;
@@
Figures 1(e) and 1(f) compare the simulated and emulated spatial fields of the 50th replicate by plotting their quantiles against each other (when pooling the spatial data into the same plot). We see that both emulations align very well with the simulated data set. Although this might not be the most appropriate way of evaluating the quality of the emulations because there is spatial dependence and non-stationarity within each spatial replicate, QQ-plots still provide value in determining whether the spatial distribution is similar at all quantile levels, which is complementary to the empirical $\chi_{ij}(u)$ described in Section B.2.
@@
D.5 Stochastic gradient descent optimization
A major advantage of approximating the ELBO as presented in Eq. (D.4) lies in the ability to perform joint optimization over all parameters ($\boldsymbol{\phi}_{e}$ and $\boldsymbol{\phi}_{d}$) using stochastic gradient descent (SGD). This optimization is efficiently implemented using a tape-based automatic differentiation module called autograd within the R package torch (Falbel and Luraschi, 2023). Built on PyTorch, this package offers rapid array computation, leveraging robust GPU acceleration for enhanced computational efficiency. It stores all the data inputs and VAE parameters in the form of torch tensors, which are similar to R multi-dimensional arrays but are designated for fast and scalable matrix calculations and differentiation.
@@ D.5 Stochastic gradient descent optimization
Importantly, our XVAE algorithm can scale efficiently to massive spatial data sets. The existing max-stable, inverted-max-stable, and other spatial extremes models are limited to applications with less than approximately $1,000$ locations using a full likelihood or Bayesian approach; see Section 2.1 of the main paper for more details on these alternative approaches. By contrast, our approach can fit a globally non-stationary spatial extremes process, with parameters evolving over time, to a data set of unprecedented spatial dimension of more than $16,000$ locations, and also facilitates data emulation in such dimensions. See Section 5 of the main paper for details.
@@
$(\boldsymbol{\phi}_{e}^{(j)},\boldsymbol{\phi}_{d}^{(j)})^{\top}\leftarrow$ Initialized parameters;
@@
             $(\boldsymbol{\mu}_{t}^{\top},\log\boldsymbol{\zeta}_{t}^{\top})^{\top}\leftarrow\mathrm{EncoderNeuralNet}_{\boldsymbol{\phi}_{e}^{(j)}}(\boldsymbol{x}_{t})$;
@@
             $(\alpha_{t},\boldsymbol{\gamma}_{t}^{\top})^{\top}\leftarrow\mathrm{DecoderNeuralNet}_{\boldsymbol{\phi}_{d}^{(j)}}(\boldsymbol{z}_{t})$;
@@
       Apply update: $(\boldsymbol{\phi}_{e}^{(j+1)},\boldsymbol{\phi}_{d}^{(j+1)})^{\top}\leftarrow(\boldsymbol{\phi}_{e}^{(j)},\boldsymbol{\phi}_{d}^{(j)})^{\top}+\boldsymbol{v}$;
       $\boldsymbol{L}\leftarrow(\boldsymbol{L}^{\top},\mathcal{L}_{\boldsymbol{\phi}_{e}^{(j)},\boldsymbol{\phi}_{d}^{(j)}}(\mathcal{M}))^{\top}$ ;
@@
Algorithm 2 Stochastic Gradient Descent with momentum to maximize the ELBO defined in Eq. (D.4). We set $|\mathcal{M}|=n_{t}$ and $L=1$ in our experiments.
D.6 Finding starting values
@@
$\hat{\boldsymbol{z}}_{t}\approx\{(\boldsymbol{\Omega}^{\frac{1}{\alpha}})^{\top}\boldsymbol{\Omega}^{\frac{1}{\alpha}}\}^{-1}(\boldsymbol{\Omega}^{\frac{1}{\alpha}})^{\top}\boldsymbol{x}_{t}^{\frac{1}{\alpha_{0}}},\;t=1,\ldots,n_{t}.$
We use QR decomposition to solve the following linear system to get the initial value $\boldsymbol{W}_{1}^{(0)}$: $(\hat{\boldsymbol{z}}_{1},\cdots,\hat{\boldsymbol{z}}_{n_{t}})^{\top}=(\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n_{t}})^{\top}\boldsymbol{W}_{1}^{\top}$. Also, set $\boldsymbol{b}_{1}^{(0)}=(0,\ldots,0)^{\top}$. The initial values of $\boldsymbol{h}_{1,t}$ in Eq. (D.1) satisfy $\boldsymbol{h}_{1,t}\approx\hat{\boldsymbol{z}}_{t}$, $t=1,\ldots,n_{t}$.
@@
Appendix E Additional results from the simulation study
We show additional figures that are complementary to those included in Section 4 of the main paper. Figure E.1 displays the simulated data sets from Models I, II, IV and V and their emulated fields using both XVAE and hetGP. See Figure 3 of the main paper for comparison for Model III. Figure E.2 displays QQ-plots from the spatial data to compare the overall distributions of the simulated and emulated data sets. Figure E.3 compares the empirically estimated $\chi_{h}(u)$ as described in Section B.2 from the data replicates simulated from Models I, II, IV and V and their emulations at three different distances $h=0.5,2,5$ under the working assumption of stationarity. Figure E.4 shows the estimates of $\mathrm{ARE}_{\psi}(u)$ defined in Eq. (12) of the main paper, $\psi=0.05$, for both simulations and XVAE emulations under Models I, II, IV and V. See Figure 5 of the main paper for $\chi_{h}(u)$ and $\mathrm{ARE}_{\psi}(u)$ estimates for Model III. Lastly, Figure E.5 shows coverage probabilities of $\{\gamma_{kt}:k=1,\ldots,K\}$ for $t=1$ from fitting Model III. Coverage probabilities when $\gamma_{k}=0$ are poor, though upper bounds of credible intervals are consistently less than $10^{-6}$.
Figure E.1: Simulated data sets (left column) and emulated fields (XVAE, middle column; hetGP, right column) from Models I, II, IV and V (top to bottom). In all cases, we use data-driven knots for emulation using XVAE. See Figure 3 of the main paper for comparison for Model III.
@@ Appendix E Additional results from the simulation study
Figure E.3: The empirically-estimated tail dependence measure $\chi_{h}(u)$ at $h=0.5$ (left), 2 (middle), 5 (right) for Models I, II, IV and V (top to bottom), based on simulated (black) and XVAE emulated (red) data. See Figure 5 of the main paper for $\chi_{h}(u)$ estimates for Model III.
Figure E.4: Estimates of $\mathrm{ARE}_{\psi}(u)$, $\psi=0.05$, for both simulations (black) and XVAE emulations (red) under Models I, II, IV and V (left to right). See the right panel of Figure 5 of the main paper for $\mathrm{ARE}_{\psi}(u)$ estimates for Model III.
@@
E.1 Nonstationary space-time dependence setting
Here, we simulate data based on the model setting III, but we additionally impose a single linear time trend to all knots. We generate 100 time points ($n_{t}=100$) at the same 2,000 spatial locations ($n_{s}=2,000$) as described in Section 4.1 of the main paper. To evaluate the model’s ability to capture temporal nonstationarity, we train the XVAE on the true knots and use the trained decoder to generate 1,000 samples to estimate $(\alpha_{t},\boldsymbol{\gamma}_{t}^{\rm T})^{\rm T}$. Figure E.6 presents the median estimates of $\{\gamma_{kt}:k=1,\dots,K,t=1,\dots,n_{t}\}$, averaged over time (left panel) and space (right panel).
The results show that our method effectively captures temporal nonstationarity, though some temporal stochastic fluctuations appear due to the working independence assumption of the max-id model and the natural variability of the estimator. This highlights an area where a conditional VAE could be particularly useful, as it could allow the encoder and decoder parameters to vary with time and other conditioning variables. This flexibility would introduce temporal change in dependence structure through the time-varying tilting parameters and enable the modeling of more complex, nonlinear temporal trends. On the spatial side, our method continues to emulate the variation in the tilting parameters well, although the estimates near the edges of the spatial domain are slightly less accurate. In comparison, hetGP does not naturally handle non-stationarity over space and time.
Figure E.6: Left: Time-averaged relative differences between the estimated and true tilting parameters at $K=25$ true knots (i.e., $n_{t}^{-1}\sum_{t=1}^{n_{t}}(\hat{\gamma}_{kt}-\gamma_{kt})/\gamma_{kt}$, $k=1,\ldots,K$). Right: Estimated tilting parameters averaged over space (i.e., $K^{-1}\sum_{k=1}^{K}\hat{\gamma}_{kt}$, $t=1,\ldots,n_{t}$), overlaid with the true trend (blue line) and the best median regression fit (red line).
E.2 Comparison to extGAN
The extGAN proposed by Boulaguiem et al. (2022) uses convolutional neural networks (CNNs) in both its generator and discriminator, constraining the spatial input to a regular grid. Consequently, for a fair comparison of the emulation performance between our XVAE and extGAN, we must use their specific simulations setup, including the same grid size and number of spatial locations. Altering the number of locations would require a complete re-design and tuning of the neural network architecture to accommodate the new dimensions.
Figure E.7: The left panel presents knot locations used for Model III but a different spatial setting to be consistent with extGAN, and we only show the support of the one Wendland basis function centered at knot in the middle of the domain. The right panels display the $\gamma_{k}$ values, $k=1,\ldots,K$, used in the expPS variables. The circled knots signify $\gamma_{k}=0$, which induces local AD.
Here, we simulate from the max-id model setting (i.e., Model III from Section 4) with a non-stationary dependence structure. Specifically, we use an $18\times 22$ grid, with $K=24$ evenly-spaced knots, employing compactly supported Wendland basis functions centered at each knot with $r=6$. As in Section 4, we use a mix of positive and zero $\gamma_{k}$’s; see Figure E.7 for an illustration of the knot placement and $\{\gamma_{k}\}$ values. Since extGAN assumes stationary marginal distributions at each grid site, we only consider time-invariant dependence parameters $\alpha_{t}\equiv 1/2$ and $\boldsymbol{\gamma}_{t}\equiv\boldsymbol{\gamma}$. Following the simulation setup for precipitation and temperature applications in Boulaguiem et al. (2022), we simulate $n.t=500$ independent replicates.
The extGAN implementation by Boulaguiem et al. (2022) was coded in TensorFlow version 1.0, which is incompatible with Python versions $\geq$ 3.0 and modules such as pandas and tensorflow_probability. Additionally, it includes inconsistencies with TensorFlow operations in Keras layers, hindering the execution of tf.function compilation. To address these limitations, we translated their GAN implementation to a tf.keras.Model class; this modified code is available in our GitHub repository at https://github.com/likun-stat/XVAE.
In general, GANs aim to generate diverse images that resemble the overall distribution of the training data rather than replicate any one specific image. As outlined in Algorithm 1 of Boulaguiem et al. (2022), the extGAN is indeed trained to generate random images from the empirical copula—the empirical joint distribution of the data after a rank transformation. Figure E.8 illustrates three rank-transformed inputs (from the empirical copula) used to train extGAN (top) and three generated images on the copula scale (bottom), showing that extGAN seeks to capture the copula’s overall dependence structure without emulating single realizations. Although GAN inversion or generator overfitting could enable emulation of specific training images, this approach diverges from standard GAN usage.
Figure E.8: Top panels show the copulas (marginally transformed to standard uniform distributions using rank transformation) from the first three time replicates. Note Boulaguiem et al. (2022) padded the copulas with 0’s at the periphery of the domain for the application of CNNs. Bottom panels show three copulas generated by extGAN, in which the GAN does not try to replicate a specific image but rather the overall patterns of the spatial distribution.
Also, due to the use of CNNs, extGAN is not able to handle missing locations across space, preventing out-of-sample emulation performances assessment via CRPS. Instead, we compare the overall dependence structures from both emulation methods to the original simulated data. Figure E.9 compares the overall marginal distribution obtained by pooling data across space and time. The emulated copula from extGAN is transformed to the data scale using the analytic marginal distribution function derived in Proposition 3.1 with the true parameters. Note that in real data applications, we have to design the marginal models carefully and estimate the model parameters well if we want to use extGAN. The right panel of Figure E.9 shows that, in this case, the marginal data quantiles are quite severely underestimated when using extGAN emulations even though we used the true parameters. This is largely due to the mis-characterization of the dependence structure in the copula. The left panel of Figure E.10 shows the comparison of MSPE in log scale (see definition in Section B.1), confirming that the quality of emulation is a bit higher when a parametric constraint on the copula is imposed, as with our XVAE. Furthermore, we compare the ARE estimates from the simulated data and the emulations. The right panel of Figure E.10 shows that the length scale of threshold exceedance is slightly overestimated for large quantile levels, although the corresponding uncertainty bands are a bit wider and contain the estimates from the simulated data (i.e., the “truth”) and the XVAE emulations.
Figure E.9: QQ-plots comparing simulated data sets and emulated fields from XVAE (left) and extGAN (right) based on Model III. Data are pooled across space and time.
For this simulated dataset, extGAN also takes longer to train ($\sim$3 hours) compared to our XVAE ($\sim$ 30 minutes). In addition, in real data applications where the marginal parameters are time-varying, generating arbitrary random images (even on the copula scale) may be not sensible, especially for the purpose of emulation.
Figure E.10: On the left, we show the MSPE values from two emulation approaches on the dataset simulated from Model III. Lower MSPE values indicate better emulation results. On the right, we show $\mathrm{ARE}_{\psi}(u)$ with $\psi=1$ based on data replicates (black), XVAE emulated data (red) and extGAN emulated data (blue).